# Surveying the Landscape of Image Captioning Evaluation: A Comprehensive Taxonomy and Novel Ensemble Method

### v1

@@ Abstract
The task of image captioning has recently been gaining popularity, and with it the complex task of evaluating the quality of image captioning models. In this work, we present the first survey and taxonomy of over 70 different image captioning metrics and their usage in hundreds of papers. We find that despite the diversity of proposed metrics, the vast majority of studies rely on only five popular metrics, which we show to be weakly correlated with human judgements. Instead, we propose EnsembEval — an ensemble of evaluation methods achieving the highest reported correlation with human judgements across 5 image captioning datasets, showing there is a lot of room for improvement by leveraging a diverse set of metrics. [^1] Our code and data are available on
@@ 1 Introduction
Image captioning is a multi-modal task in which models generate textual captions describing an image. Evaluating the output of image captioning is challenging for several reasons. First, as in other generative tasks, multiple outputs may be valid for the same input, as illustrated in Figure 1, where different valid captions for the same image have no overlap in content words. Second, it involves bridging across modalities, requiring evaluators to compare text and images, unlike most generative tasks that only involve textual information. Attesting to the difficulty and importance of image captioning is the sheer volume of metrics proposed for this task, and their usage in hundreds of image captioning models (Figure 2).
In this work, we survey the different approaches proposed for evaluating image captioning, and provide a first taxonomy explaining over 70 metrics, which were largely developed independently of one another. Furthermore, based on our observations we devise an ensemble-based metric, EnsembEval, which achieves the highest reported correlation with human ratings.
@@
Figure 2: Metrics usage over the years. Metrics used in less than two papers are omitted.
@@
In Section 2, we start our survey with a systematic approach to examine all relevant papers from 15 major venues in NLP, vision and machine learning between 2010 and 2024. This approach yields a body of work consisting of 71 different automatic evaluation metrics and 5 human evaluation paradigms, used in over 300 papers.
@@
Next, we organize both automatic metrics and human evaluation in principled taxonomies (Section 3). Our taxonomy is the first to comprehensively cover all automatic metrics used in image captioning research, organized by a novel criterion: the specific property each metric aims to quantify. This enables future researchers to choose metrics based on the specific aspect their model is intended to improve. In addition, we propose the first taxonomy for human evaluation metrics, categorizing them into groups such as comparative evaluation and scale rating evaluation.
@@
We then survey the use of evaluation methods across a wide range of 314 papers (Section 4). Interestingly, we find that despite the wealth of different metrics proposed for the task, the vast majority of examined papers use only five simple metrics (BLEU, METEOR, ROUGE, CIDEr, SPICE), although there has been a the recent increase in the adoption of alternative metrics. Furthermore, we find that the use of human evaluation is declining in recent years and that significance and inter-annotator agreement are rarely reported.
@@
Finally, in a series of evaluations, we show that the five most popular metrics show only a weak correlation with human ratings, whereas lesser-known metrics show a much higher correlation. This underscores the value of a systematic and comprehensive survey in identifying these lesser-known but more effective metrics. We then propose an ensemble of selected metrics, EnsembEval, optimized for diversity using a feature selection algorithm. We show that EnsembEval achieves the highest reported correlation with human ratings to date.
@@
To recap, our contributions are twofold. First, we introduce the first comprehensive taxonomy of automatic captioning evaluation metrics, categorized by the property they assess. This will help future model developers select the most suitable metric based on the specific property they want their model to excel in. Second, our proposed ensemble approach based on this survey can be readily used as a state-of-the-art evaluation of image captioning models.
2 Methodology
@@
We perform a systematic review and gather usage statistics of metrics in conference papers on English image captioning. Overall, we examined 314 papers from 2010 to 2024, and identified 71 distinct metrics.
@@
Venues.
We focus on a large set of 15 venues spanning three distinct research communities (Table 1). Figure 3 shows the annual number of papers for each community.
@@
Search strategy.
Filtering.
Next, we manually filter out papers that we deem irrelevant to our review. Specifically, we exclude: 1) papers on other types of captioning such as video captioning; 2) papers unrelated to captioning that happen to have caption or description in their titles; 3) papers that focus on languages other than English; and 4) papers that did not conduct any experiments.
Years reviewed.
We begin collecting data starting from 2024 backwards until reaching the first year with no relevant papers (2009).
Collected data.
@@ 3 Taxonomy
In this section, we introduce a taxonomy of the metrics from the systematic review outlined in Section 2. This work is novel in two respects: we are the first to comprehensively include all metrics proposed and used in previous studies, rather than focusing on a limited set of well-known metrics. Second, we are the first to describe a taxonomy for human evaluation methods.
@@ 3.1 Automatic Evaluation
Figure 4: A taxonomy of automatic metrics for image evaluation. For each category, we indicate the number of metrics within that category in parentheses. Top categories are preceded by the relevant section number. Categories are in blue, metrics in red. * – Nubia is assigned to two distinct categories under Relation with ground-truth.
The release of large benchmarks like MSCOCO Lin et al. (2014) has made manual evaluation of captioning models impractical. Hence, automatic tools for assessing generated captions have been developed, and some tools from other text generation tasks have been adapted for the image captioning task.
We turn to describing the various automatic evaluation methods. We categorize metrics based on the specific property they aim to evaluate. [^3] One metric (Yngve score, Liu et al., 2019a) could not fit into any of the proposed categories, while another (Nubia) is assigned to two categories. The taxonomy is presented visually in Figure 4.
@@ 3.2 Human Evaluation
While automatic evaluation offers clear benefits in terms of scale and consistency, it still serves as a surrogate for human evaluation. To ensure that the improvements shown by automatic methods are genuine, many image captioning studies also apply human evaluation methods to a subset of the data. Here, we categorize various frameworks for human evaluation.
@@ 4.1 Automatic Evaluation
Figure 2 displays the number of papers using different metrics each year. Given the large number of metrics identified (71), we only plot usage for the seven most common metrics, clustering all other metrics by the categories from Section 3. Metrics used by no more than one paper are omitted.
The figure shows that since 2015, five metrics – BLEU, CIDEr, METEOR, ROUGE, and SPICE (henceforth, the five dominant metrics) – have been used substantially more frequently than all other metrics. This trend has begun to shift in recent years, with an increase in the use of other metrics starting from 2020. Notably, four of the dominant metrics (all except SPICE) are lexical similarity metrics, criticized for failing to capture semantic similarity Giménez and Màrquez (2007), and none consider the image during evaluation.
Figure 5: Metrics usage analysis plots. Top: Mean number of metrics used per paper each year. Center: Fraction of papers that performed human evaluation each year. Bottom: Human evaluation frameworks usage over the years, smoothed by convolving with a window size of four years.
@@
Figure 5 (top) shows the mean number of metrics used per paper each year. There is a gradual increase until the mean number reached approximately 4.5 in 2019, remaining in the range of 4-5 since then.
@@ 4.2 Human Evaluation
Figure 5 (center) shows the fraction of papers with human evaluation each year. From 2010 to 2014, all papers conducted human evaluation. However, the publication of large image-caption datasets made this more challenging, and beginning at 2015 (a year after MSCOCO’s release) there has been a general decline in the use of human evaluation, likely due to improvements in automatic evaluation metrics.
@@
Figure 5 (bottom) presents the number of papers using each human evaluation framework per year. The comparative framework is most frequently used, closely followed by scale rating.
@@
Out of the 94 papers that used human evaluation (29.9% of the papers we documented), only 19 reported significance. Reporting significance is crucial in human evaluation because, in most cases, only a small subset of the test set is assessed due to the high cost of human labor. Moreover, the results may be unreliable if the annotator agreement is low; however, only 6 of these 94 papers reported annotator agreement.
@@ 5 Experiments
Following previous work Anderson et al. (2016); Hessel et al. (2021), we use correlation with human ratings as an indicator of metric quality. While inspired by previous work, our experiments include metrics not studied previously.
@@ 5.1 Experimental Setup
THumB Kasai et al. (2022) features human ratings of captions for 500 images sourced from MSCOCO. Each image is associated with four model-generated and one original reference captions. Ratings cover various dimensions, such as precision, recall, and an overall score. We take the overall score as the score for this caption.
Polaris Wada et al. (2024) contains 131,020 human ratings for captions generated for 13,691 images from MSCOCO and nocaps Agrawal et al. (2019). The captions were generated by 10 different models. The dataset is partitioned into training, validation and test sets.
@@
The Reformulations dataset (introduced in a work currently under review) includes model-generated captions for 1,405 images from MSCOCO and Flickr30k, along with human-generated reformulations, i.e., corrected versions of the captions (if any errors exist). Evaluation metrics are expected to favor the caption after human correction. The dataset contains 5,208 caption-reformulation pairs; we omit 864 pairs where the reformulation is identical to the caption.
@@
Metrics selection.
We now select the metrics to be evaluated out of the 71 identified metrics. First, we exclude bias metrics, as bias is not captured by the human ratings datasets. Next, we filter out all metrics that receive multiple candidates as input (such as diversity metrics), as the human ratings are provided for a single candidate. Finally, we exclude all metrics that rely on a closed model API access (CLAIR, GPT4V evaluation, ALOHa), focusing on publicly accessible evaluation methods.
After filtering, 56 metrics remain. To ensure fair comparison, we conduct all the experiments for each metric ourselves, even if previous studies have reported results for this metric. Due to the effort required and the scope of this study, we focus on metrics frequently employed in image captioning research. Specifically, we select metrics used in at least two papers per year on average since their publication. [^4] Not including the publication year and 2024. This narrows our selection to 20 metrics. [^5] BLEU1, BLEU2, BLEU3, BLEU4, BLIP2Score, CIDEr, CLIPImageScore, CLIPScore, Exact noun overlap (NO), Exact verb overlap (VO), Fuzzy noun overlap (NO), Fuzzy verb overlap (VO), METEOR, MPNetScore, PACScore, ROUGE, RefCLIPScore, RefPACScore, SPICE, Polos. We omit two metrics for which we could not obtain a full implementation: PR-MCS and InfoMetIC. For details on the implementation of each metric, refer to Appendix A.2.
@@ 5.2 EnsembEval: An Ensemble of Evaluation Methods
Metric Category Coef
BLIP2Score Image similarity 0.83
Polos Human rating prediction 0.82
PACScore Image similarity 0.29
Exact NO Lexical similarity 0.08
BLEU1 Lexical similarity 0.07
Fuzzy VO Phrase semantic similarity 0.02
BLEU4 Lexical similarity 0.02
CIDEr Lexical similarity -0.02
ROUGE Lexical similarity -0.07
RefCLIPScore Multiple source similarity -0.16
Table 2: Metrics selected by the feature selection algorithm to predict human ratings, along with their categories and linear regression coefficients.
@@
Table 3: Correlation with human ratings across different datasets. The best performing metric is in bold. Metrics are sorted by their average score.
We now examine whether a more diverse evaluation method, incorporating multiple existing metrics, outperforms individual metrics. We propose an ensemble approach we name EnsembEval, using a linear combination of these metrics.
@@
Next, we find the metrics coefficients by training a linear regression model to predict human ratings, again using the Polaris train dataset. More details on our feature selection and linear regression procedures can be found in Appendix A.3.
Table 2 displays the selected metrics, their taxonomy categories, and their linear regression coefficients. The metrics span different categories, indicating that predicting human ratings could benefit from using different types of metrics. Interestingly, some metrics received a negative coefficient, most notably RefCLIPScore which is widely regarded as highly correlated with human ratings Kasai et al. (2022).
@@
Table 4: Accuracy in pairwise comparison. In Pascal50S we report mean and standard deviation across five random instances of tie-breaking and reference selection. HC: both captions are human-generated for the image, HI: both are human-generated but only one corresponds to the image, HM: one caption is human-generated while the other is machine-generated, MM: both captions are machine-generated. In each dataset, best scoring metric is in bold. Metrics are sorted by their average score.

### v2

# Surveying the Landscape of Image Captioning Evaluation: A Comprehensive Taxonomy, Trends and Metrics Analysis

@@ Abstract
The task of image captioning has recently been gaining popularity, and with it the complex task of evaluating the quality of image captioning models. In this work, we present the first survey and taxonomy of over 70 different image captioning metrics and their usage in hundreds of papers, specifically designed to help users select the most suitable metric for their needs. We find that despite the diversity of proposed metrics, the vast majority of studies rely on only five popular metrics, which we show to be weakly correlated with human ratings. We hypothesize that combining a diverse set of metrics can enhance correlation with human ratings. As an initial step, we demonstrate that a linear regression-based ensemble method, which we call EnsembEval, trained on one human ratings dataset, achieves improved correlation across five additional datasets, showing there is a lot of room for improvement by leveraging a diverse set of metrics. [^1] Our code and data are available on
@@ 1 Introduction
Evaluating the output of image captioning is challenging for several reasons. First, as in other generative tasks, multiple outputs may be valid for the same input, as illustrated in Figure 1, where different valid captions for the same image have no overlap in content words. Second, it involves bridging across modalities, requiring evaluators to compare text and images, unlike most generative tasks that only involve textual information. Attesting to the difficulty and importance of image captioning is the sheer volume of metrics proposed for this task, and their usage in hundreds of image captioning models (Figure 2).
In this work, we survey the different approaches proposed for evaluating image captioning, and provide a first taxonomy comprising over 70 metrics, which were largely developed independently of one another. We examine usage patterns of metrics over the past 14 years and demonstrate that combining multiple metrics enhances correlation with human ratings.
@@
Figure 2: Metrics usage over the years. Metrics used in $\leq 1$ papers are omitted. $N=$ indicates number of metrics in this category.
@@
We begin by exploring various definitions of image captioning presented in previous studies and examining their impact on evaluation (Section 2). Then, in Section 3, we systematically examine all relevant papers from 15 major venues in NLP, vision and machine learning between 2010 and 2024. This approach yields a body of work consisting of 71 different automatic evaluation metrics and 5 human evaluation paradigms, used in over 300 papers.
@@
Next, we organize both automatic metrics and human evaluation in a principled taxonomy (Section 4). Our taxonomy is the first to comprehensively cover all automatic metrics used in image captioning research. We design the taxonomy dimensions with the needs of metric users in mind, focusing on dimensions such as the candidate’s property each metric aims to measure. We also propose the first taxonomy for human evaluation metrics, categorized into groups such as comparative evaluation and scale rating evaluation.
@@
We then survey the use of evaluation methods across 314 papers (Section 5). Interestingly, we find that despite the wealth of different metrics proposed for the task, the vast majority of examined papers use only five simple metrics (BLEU, METEOR, ROUGE, CIDEr, SPICE), although there has been a recent increase in the adoption of alternative metrics. Furthermore, we find that the use of human evaluation is declining in recent years and that significance and inter-annotator agreement are rarely reported.
@@
Finally, in a series of evaluations, we show that the five most popular metrics show only a weak correlation with human ratings, whereas lesser-known metrics show a much higher correlation. This underscores the value of a systematic and comprehensive survey in identifying these lesser-known but more effective metrics. We then show that an ensemble of selected metrics we name EnsembEval, optimized for diversity using a feature selection algorithm, achieves enhanced correlation with human ratings. This paves the way for future research on metrics that consider multiple aspects simultaneously.
@@
To recap, our contributions are threefold. First, we introduce the first comprehensive taxonomy of automatic captioning evaluation metrics categorized by user-oriented aspects, helping future model developers select the most suitable metric based on the specific property they want their model to excel in. Second, we examine key trends in image captioning evaluation, highlighting a shift from reliance on five simple metrics toward alternative approaches. Finally, we demonstrate that a simple linear regression-based ensemble method improves correlation with human ratings and provide the code publicly for future use.
2 Task Definition
Image captioning can be motivated by diverse goals, and what counts as a ‘good’ caption heavily depends on the system’s purpose and context. Accurate evaluation requires explicitly defining a system’s purpose, a step often overlooked or only implicitly addressed in previous work. We start with a review of common definitions.
Early research framed image captioning as a means to assess vision-and-language models based on image-caption semantics rather than practical application Hodosh et al. (2013). Recently the task is commonly defined as describing the visual content of an image or scene in natural language Stefanini et al. (2022); Nallapaneni and Konakanchi (2023); González-Chávez et al. (2024). Yet, this definition does not specify what visual content should be described or the intended nature of the description. These aspects depend on both the purpose and the context of the description, which we now discuss in detail.
Descriptions serve different purposes, each with distinct requirements. For example, captions designed for accessibility should replace the image, while captions in news articles typically augment or link the content of the image and article Kreiss et al. (2022). In practice, although often implicit, most previous studies focus on image replacement, i.e., generating descriptions intended to serve as substitutes for the image, as defined by Kreiss et al. (2022).
The second key factor is the context in which a caption is generated. For instance, the top caption in Figure 1 may suit Lego’s Instagram account, while the center caption might be more appropriate for a travel guide.
While our survey covers all works on captioning evaluation – irrespective of purpose or context – our experiments on metric performance and ensembling (Section 6) focus on the replacement task, specifically .
3 Literature Survey
@@
We perform a systematic review and gather usage statistics of metrics in conference papers on English image captioning. Overall, we examined 314 papers, and identified 71 distinct metrics.
@@
Scope.
We focus on a large set of 15 venues spanning three distinct research communities (Table 1). Figure 3 shows the annual number of papers for each community. We begin collecting data starting from 2024 backwards until reaching the first year with no relevant papers, which is 2009.
Search strategy.
Filtering.
Next, we manually filter out papers that we deem irrelevant to our review. Specifically, we exclude: 1) papers on other types of captioning such as video captioning; 2) papers unrelated to captioning that happen to have caption or description in their titles; 3) papers that focus on languages other than English; and 4) papers that did not conduct any experiments (e.g., review papers).
Collected data.
@@ 4 Taxonomy
Here, we introduce a taxonomy of the metrics from the systematic review outlined in Section 3.
This work is novel in two respects: we are the first to comprehensively include all metrics proposed and used in previous studies, rather than focusing on a limited set of well-known metrics. Second, we are the first to describe a taxonomy for human evaluation methods.
@@ 4.1 Automatic Evaluation
Figure 4: A taxonomy of automatic metrics for image captioning evaluation. For each category, we list the number of metrics in parentheses. Top categories are preceded by the relevant section number. Blue rectangles contain categories, red contain metrics. * – Nubia is assigned to two distinct categories under Relation with ground-truth.
Motivation.
Given the large number of 71 automatic evaluation methods identified in our survey, it is a challenge to stay updated on the latest metrics. As a result, the common practice is to rely on previously used metrics, typically the five dominant ones – BLEU, CIDEr, METEOR, ROUGE, and SPICE – as shown in Figure 2.
To address this issue, we propose a taxonomy that helps users select the most suitable metric. We categorize metrics based on the specific property they assess in the candidate caption.
Method.
We begin by categorizing the metrics based on the evaluated property. For example, BLEU, BERTScore and CLIPScore are classified under lexical similarity to reference, phrase level semantic similarity to reference and semantic similarity to image, respectively. [^3] One metric (Yngve score Liu et al. (2019a)) did not align with any category, another (Nubia Kane et al. (2020)) fit into two categories.
Next, we incrementally group together pairs of categories based on their similarity to one another. For instance, lexical similarity to reference and phrase level semantic similarity to reference are both grouped under similarity to reference, while semantic similarity to image falls under similarity to image. We repeat this process iteratively until only one overarching category remains.
Finally, following the process of Nickerson et al. (2013), we extend the taxonomy with three additional dimensions, again motivated by the overarching goal to support users in their choice of metrics:
  • Input type: The data provided to the metric (candidate only, candidate + reference, candidate + image, candidate + reference + image).
  • Implementation availability: Availability of the metric’s implementation (none, dedicated codebase, or integration within a generic platform like HuggingFace).
  • API access requirements: Whether the metric relies on a closed model API access.
The initial taxonomy was constructed by the first author, followed by validation from the remaining authors. Any disagreements were resolved through majority voting. The final taxonomy is visually presented in Figure 4. In the following sections, we describe the various automatic evaluation methods.
@@ 4.2 Human Evaluation
While automatic evaluation offers clear benefits in terms of scale and consistency, it still serves as a surrogate for human evaluation. To ensure that the improvements shown by automatic methods are genuine, many image captioning studies also apply human evaluation methods to a subset of the data.
The human evaluation taxonomy development process followed a similar approach to that in automatic evaluation (Section 4.1). In this case, a single dimension emerged: the evaluation framework.
@@ 5.1 Automatic Evaluation
Figure 2 displays the number of papers using different metrics each year. Given the large number of metrics identified (71), we only plot usage for the seven most common metrics, clustering all other metrics by the categories from Section 4.
The figure shows that since 2015, five metrics – BLEU, CIDEr, METEOR, ROUGE, and SPICE (henceforth, the five dominant metrics) – have been used substantially more frequently than all other metrics. This trend has begun to shift in recent years, with an increase in the use of other metrics. Notably, four of the dominant metrics (all except SPICE) are lexical similarity metrics, criticized for failing to capture semantic similarity Giménez and Màrquez (2007). Figure 5 (A) illustrates a decline in the usage of lexical metrics in recent years relative to overall metric usage, excluding diversity and bias metrics from this count.
Figure 5: Metrics usage analysis plots. (A): Percentage of lexical metric usage relative to total metric usage each year, smoothed by convolving with a window size of four years. (B): Mean number of metrics used per paper each year. (C): Fraction of papers that performed human evaluation each year. (D): Human evaluation frameworks usage over the years, smoothed by convolving with a window size of four years.
@@
Figure 5 (B) shows a gradual increase in the mean number of metrics used per paper each year until reaching approximately 4.5 in 2019, and remaining in the range of 4-5 since then.
@@ 5.2 Human Evaluation
Figure 5 (C) shows the fraction of papers with human evaluation each year. From 2010 to 2014, all papers conducted human evaluation. However, the publication of large image-caption datasets made this more challenging, and beginning at 2015 (a year after MSCOCO’s release) we see a general decline in the use of human evaluation, likely due to improvements in automatic evaluation metrics.
@@
Figure 5 (D) presents the number of papers using each human evaluation framework per year. The comparative framework is most frequently used, closely followed by scale rating.
@@
Of 94 papers that used human evaluation (29.9% of the papers we documented), only 19 reported significance, which is crucial in human evaluation because, due to the high cost of human labor, only a small subset of the test set is assessed. Moreover, results may be unreliable if annotator’s agreement is low; however, only 6 of these 94 papers report annotator agreement.
@@ 6 Experiments
Prior captioning evaluation work assumed an automatic metric’s quality depends on its alignment with human evaluations. To facilitate this, human rating datasets were created, where participants rated candidate captions. A common approach (among others, Anderson et al., 2016; Hessel et al., 2021) is to evaluate metrics by applying them to the candidates in these datasets and measuring the correlation of their scores with the human ratings. We follow this convention but extend prior research by incorporating metrics that have not been previously studied in our experiments.
Throughout our experiments, we focus on the replacement subtask (see Section 2). This definition – producing descriptions intended to serve as substitutes for the image – aligns most closely with the human rating datasets, where annotators are instructed to evaluate whether the text accurately reflects the image’s content without adding information not present in the image. [^4] See Appendix A.1 for annotators instructions for existing human rating datasets.
@@ 6.1 Experimental Setup
THumB Kasai et al. (2022) features human ratings of captions for 500 images sourced from MSCOCO. Each image is associated with four model-generated and one original reference captions. Annotators provide ratings for precision and recall (on a scale of 1 to 5), deducting some penalty points from their average in case of lack of fluency, conciseness and incluse language to reach the overall score. We take this overall score as the score for this caption.
Polaris Wada et al. (2024) contains 131,020 human ratings for captions generated for 13,691 images from MSCOCO and nocaps Agrawal et al. (2019). Ratings were provided on a five-point scale and then transformed to values in the range [0,1] using min-max normalization. The captions were generated by 10 different models. The dataset is partitioned into training, validation and test sets.
@@
Cluster Metrics
BLEU1, BLEU2, BLEU3, BLEU4,
1 CIDEr, Exact NO, Fuzzy NO,
METEOR, ROUGE, SPICE
BLIP2Score, CLIPImageScore,
2 CLIPScore, MPNetScore, PACScore,
RefCLIPScore, RefPACScore, Polos
3 Exact VO, Fuzzy VO
Table 2: Metrics clustered by mutual correlation.
The Reformulations dataset Berger et al. (2025) includes model-generated captions for 1,405 images from MSCOCO and Flickr30k, along with human-generated reformulations, i.e., corrected versions of the captions (if any errors exist). Evaluation metrics are expected to favor the caption after human correction. The dataset contains 5,208 caption-reformulation pairs; we omit 864 pairs where the reformulation is identical to the caption.
@@ 3 Exact VO, Fuzzy VO
Metrics.
We select a subset of the 71 identified automatic metrics for evaluation. First, we exclude bias metrics, as bias is not captured by the human rating datasets. Next, we filter out all metrics that receive multiple candidates as input (such as diversity metrics), as the human ratings are provided for a single candidate. Finally, we exclude all metrics that rely on a closed model API access (CLAIR, GPT4V evaluation, ALOHa), focusing on publicly accessible evaluation methods.
After filtering, 56 metrics remain. To ensure fair comparison, we conduct all the experiments for each metric ourselves, even if previous studies have reported results for this metric. Due to the effort required and the scope of this study, we focus on metrics frequently employed in image captioning research. Specifically, we select metrics used in at least two papers per year on average since their publication. [^5] Not including the publication year and 2024. This narrows our selection to 20 metrics (Table 2). [^6] We omit two metrics for which we could not obtain a full implementation: PR-MCS and InfoMetIC. For details on the implementation of each metric, refer to Appendix A.2.
6.2 Metric Mutual Correlation
To examine relationships among the selected metrics, we analyze their mutual correlations using Spectral Clustering with $N=3$ clusters, [^7] The number of clusters was chosen to maximize the Silhouette score. where the adjacency between two metrics is defined as their mutual correlation on the Polaris train set. The resulting grouping is shown in Table 2. Cluster 1 primarily includes lexical similarity metrics, cluster 2 consists of Transformer-based metrics, and cluster 3 contains verb overlap metrics, which exhibit low correlation with all other metrics.
@@ 6.3 EnsembEval: An Ensemble of Evaluation Methods
Metric Cluster Coef
Polos 2 0.55
BLIP2Score 2 0.40
PACScore 2 0.29
Exact NO 1 0.08
BLEU1 1 0.07
Fuzzy VO 3 0.02
ROUGE 1 -0.07
CIDEr 1 -0.15
RefCLIPScore 2 -0.17
Table 3: Metrics selected using feature selection to predict human ratings, their cluster numbers (as shown in Table 2), and linear regression coefficients.
@@
Table 4: Correlation with human ratings across different datasets. The best performing metric is in bold.
Our survey reveals that while a large number of metrics (71, as detailed in our literature review, Section 3) have been developed to assess diverse properties (outlined in our taxonomy, Section 4), only a small subset from a few categories is widely used (as shown in our usage analysis, Section 5). Furthermore, popular metric groups have faced increasing criticism in recent years. For instance, lexical similarity metrics are often criticized for failing to capture semantic similarity effectively  Giménez and Màrquez (2007), while the recently popularized reference-free metrics have been challenged for focusing solely on visual grounding errors while neglecting caption implausibility errors  Ahmadi and Agrawal (2024).
Consequently, we hypothesize that a more diverse evaluation approach, capable of capturing multiple aspects of a candidate caption, could enhance correlation with human ratings. As an initial step toward this goal, we explore a straightforward implementation: an ensemble method combining existing metrics through a linear combination. [^8] Future research can explore more sophisticated multi-aspect metrics, which lies beyond the scope of this work. We train the linear coefficients on one human rating dataset and evaluate the ensemble on additional datasets. If capturing multiple aspects indeed improves alignment with human judgments, the learned coefficients should generalize across datasets, resulting in improved correlation with human ratings. We refer to our proposed ensemble approach as EnsembEval.
@@
Next, we standardize the metrics values to the range $[0,1]$ and find the metrics coefficients by training a linear regression model to predict human ratings, again using the Polaris train dataset. The rescaling ensures that the coefficients are meaningfully comparable. [^10] Rescaling did not impact the performance of our model. More details on our feature selection and linear regression procedures can be found in Appendix A.3.
Table 3 displays the selected metrics, their cluster assignments (Section 6.2), and their linear regression coefficients. The Transformer-based metrics (cluster 2) have the largest coefficients, making them the most influential in predicting the candidate score. Negative coefficients are observed for some metrics in clusters 1 and 2, likely to mitigate redundancy from multiple metrics within the same cluster. For instance, BLIP2Score, Polos, and PACScore all receive positive coefficients but are highly correlated. To balance this redundancy, a fourth metric from the same cluster (RefCLIPScore) is assigned a negative coefficient.
@@
Table 5: Accuracy in pairwise comparison. In Pascal50S we report mean and standard deviation across five random instances of tie-breaking and reference selection. HC: both captions are human-generated for the image, HI: both are human-generated but only one corresponds to the image, HM: one caption is human-generated while the other is machine-generated, MM: both captions are machine-generated. In each dataset, best scoring metric is in bold.